# SoFlow: Solution Flow Models for One-Step Generative Modeling

### v1

@@ Abstract
Figure 1: Visual samples and quantitative comparison. (a) One-step samples generated by our Solution Flow Models on the ImageNet $\times256$ dataset. (b) With the same Diffusion Transformer (Peebles and Xie, 2023) architecture and the same number of training epochs, our models (SoFlow) consistently achieve better 1-NFE FID-50K scores than MeanFlow (Geng et al., 2025) models on the ImageNet $\times256$ dataset.
Section 1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020; Song et al., 2020) and Flow Matching models (Lipman et al., 2022; Liu et al., 2022; Albergo and Vanden-Eijnden, 2022) have emerged as foundational frameworks in generative modeling. Diffusion models operate by systematically adding noise to data and then learning a denoising process to generate high-quality samples. Flow Matching offers a more direct alternative, modeling the velocity fields that transport a simple prior distribution to a complex data distribution. Despite their power and success across various generative tasks (Esser et al., 2024; Ma et al., 2024; Polyak et al., 2024), both approaches rely on an iterative, multi-step sampling process, which hinders their generation efficiency.
Addressing this latency is becoming a key area of research. Consistency Models (Song et al., 2023; Song and Dhariwal, 2023; Geng et al., 2024; Lu and Song, 2024) and related techniques (Kim et al., 2023; Wang et al., 2025a; Frans et al., 2024; Heek et al., 2024) have gained prominence by enabling rapid, few-step generation. These methods learn a direct mapping from any point on a generative trajectory to a consistent, “clean” output, bypassing the need for iterative refinement. However, this paradigm introduces significant challenges. Consistency models trained from scratch often fail to leverage Classifier-Free Guidance (CFG) for enhancing sample quality, and they are further hindered by instability caused by changing optimization targets. Recent works (Peng et al., 2025; Geng et al., 2025) address this instability by incorporating a Flow Matching loss. However, this approach introduces a new computational bottleneck: it relies on Jacobian-vector product (JVP) calculations, which are much less optimized than forward propagation in frameworks such as PyTorch.
@@
In this work, we introduce a new approach for one-step generation that avoids these limitations. Instead of relying on iterative ODE solvers, we propose directly learning the solution function of the velocity ODE defined by Flow Matching to generate high-quality samples, as shown in Figure 1, Figure 4, and Figure 5. We denote this function as $f(x_{t},t,s)$, which explicitly maps a state $x_{t}$ at time $t$ to its evolved state $x_{s}$ at another time $s$. To learn it with a parameterized model $f_{\theta}(x_{t},t,s)$, we first analyze the properties that enable a neural network to serve as a valid solution function.
@@
Diffusion models (Sohl-Dickstein et al., 2015; Song et al., 2020; Kingma et al., 2021; Karras et al., 2022) and Flow Matching (Lipman et al., 2022; Liu et al., 2022) are widely adopted generative modeling frameworks. These approaches either progressively corrupt data with noise and train a neural network to denoise it, or learn a velocity field that governs a transformation from the data distribution to a simple prior. They have been scaled successfully for image generation (Rombach et al., 2022; Saharia et al., 2022; Podell et al., 2023) and video generation (Ho et al., 2022; Brooks et al., 2024) tasks. Stochastic interpolants (Albergo and Vanden-Eijnden, 2022; Albergo et al., 2023) build upon these concepts by explicitly defining stochastic trajectories between the data and prior distributions and aligning their associated velocity fields to enable effective distributional transport.
@@
Reducing the number of sampling steps has become an active research direction, driven by the need for faster generation and better theoretical insights. Prior efforts fall into two main streams: (i) distillation-based approaches that compress pre-trained multi-step models into few-step generators (Salimans and Ho, 2022; Sauer et al., 2024; Geng et al., 2023; Luo et al., 2023b; Yin et al., 2024; Zhou et al., 2024); and (ii) from-scratch training, which learns fast samplers without teachers, most prominently through the consistency-training family (CMs, iCT, ECT, sCT) (Song et al., 2023; Song and Dhariwal, 2023; Geng et al., 2024; Lu and Song, 2024).
Consistency Models (CMs) collapse noised inputs directly to clean data, enabling one- or few-step generation (Song et al., 2023). While first applied mainly in distillation (Luo et al., 2023a; Geng et al., 2024), later works established that they can also be trained from scratch via consistency training (Song and Dhariwal, 2023). Building on this, iCT (Song and Dhariwal, 2023) simplifies objectives and improves stability with robust losses and teacher-free training, while ECT (Geng et al., 2024) introduces a continuous-time ODE formulation that unifies CMs and diffusion. The recent sCT framework (Lu and Song, 2024) further stabilizes continuous-time consistency training, scaling CMs up to billion-parameter regimes. Beyond fixed start and end settings, newer approaches extend CMs to arbitrary timestep transitions, aligning better with continuous-time dynamics.
Specifically, Shortcut Models (Frans et al., 2024) condition on noise level and step size to flexibly support both one- and few-step sampling with shared weights. MeanFlow (Geng et al., 2025) introduces interval-averaged velocities and analyzes the relationship between averaged and instantaneous velocities. IMM (Zhou et al., 2025) matches moments across transitions in a single-stage objective, reducing variance and avoiding multi-stage distillation. Flow-Anchored CMs (Peng et al., 2025) regularize shortcut learning with a Flow Matching anchor, improving stability and generalization. On the distillation side, Align Your Flow (Sabour et al., 2025) unifies CM and FM into continuous-time flow maps effective across arbitrary step counts, further enhanced with autoguidance and adversarial fine-tuning. Finally, Transition Models (Wang et al., 2025b) reformulate generation around exact finite-interval dynamics, unifying few- and many-step regimes and achieving monotonic quality gains as step budgets increase.
Section 3 Preliminary: Flow Matching
@@
However, directly optimizing a neural network with this loss is impractical since the real velocity field ${v}({x}_{t},t)$ is a conditional expectation given ${x}_{t}$. To overcome this challenge, a conditional variant of the Flow Matching loss is introduced (Lipman et al., 2022; Liu et al., 2022):
@@
Section 4 Solution Flow Models
@@ 4.1 Formulations
The expression is equal to 0{0} due to the conditions in Equation 6 and Equation 9. Thus, we know that the model ${f}_{{{\theta}}}({x}_{t},t,s)$ is indeed the true solution function ${f}({x}_{t},t,s)$ by:
@@ 4.2 Learning Objectives
We now consider how to construct efficient objectives for neural networks to learn the ground truth solution function, according to the two conditions mentioned in the previous section. To ensure the first boundary condition in Equation 8 is satisfied, we adopt the following parameterization:
@@
where $a(t,s)$ and $b(t,s)$ are continuously differentiable scalar functions satisfying $a(t,t)=1$ and $b(t,t)=0$ for all $t\in[0,1]$, and ${F}_{{{\theta}}}({x},t,s)$ denotes a raw neural network. Following the notation in the last section, we use $\partial_{1}a(t,s),\partial_{2}a(t,s),\partial_{1}b(t,s),\partial_{2}b(t,s)$ to denote the partial derivatives of $a(t,s)$ and $b(t,s)$ with respect to the first and second variables. In our experiments, we choose two specific parameterizations, including the Euler parameterization and the trigonometric parameterization:
@@
Trivially, these two parameterizations satisfy the boundary condition ${f}_{{\theta}}({x}_{t},t,t)={x}_{t}$. Then, we construct two loss functions for training using Equation 9.
@@
Firstly, we consider a special case of Equation 9 when $t=s$. Recall that the boundary condition Equation 8 gives ${f}_{{\theta}}({x}_{t},t,t)={x}_{t},\forall{x}_{t}\in\mathbb{R}^{n},t\in[0,1]$. Thus, we have:
@@
where ${I}_{n}\in\mathbb{R}^{n\times n}$ is the identity matrix. Then Equation 9 can be simplified to ${v}({x}_{t},t)=\partial_{3}{f}_{{\theta}}({x}_{t},t,t)$. Under our parameterization mentioned above, we have
@@
Previous works (Geng et al., 2024, 2025) have demonstrated that choosing an adaptive weighting function is beneficial for few-step generative models. Following their approach, we choose
@@
The original Flow Matching framework samples $t$ uniformly, while more recent works (Esser et al., 2024; Geng et al., 2025) suggest sampling $t$ from a logit-normal distribution. We also sample $t$ from $\sigma(\mathcal{N}(\mu_{\text{FM}},\sigma_{\text{FM}}^{2}))$, where $\sigma(\cdot)$ and $\mathcal{N}$ represent the sigmoid function and normal distribution.
@@
We now consider how to build a training target for the $s<t$ situation using Equation 9. According to the Taylor expansion, we have an approximation equation:
@@
$\mathcal{L}_{\text{SCM}}({{\theta}})=\mathbb{E}_{t,l,s,{x}_{0},{x}_{1},{x}_{t}}\left[\frac{w_{\text{SCM}}(t,l,s,\text{MSE})}{n}\left\|{f}_{{{\theta}}}({x}_{t},t,s)-{f}_{{{\theta}}^{-}}\bigl({x}_{t}+(\alpha_{t}^{\prime}{x}_{0}+\beta_{t}^{\prime}{x}_{1})(l-t),l,s\bigr)\right\|_{2}^{2}\right],$ (20)
@@
where $k,K$ represent the current and total training steps, respectively. The function $r(k,K)$ represents a monotonically decreasing schedule that gradually moves $l$ towards $t$ throughout training. To avoid numerical issues, $l$ is clamped to ensure $l<t-10^{-4}$. This schedule decreases from an initial value $r_{\text{init}}$ to an end value $r_{\text{end}}$. In our ablation studies, we test exponential, cosine, linear, and constant schedules. For more implementation details, please refer to Appendix A.
@@ 4.3 Classifier-Free Guidance
Classifier-Free Guidance (CFG) (Ho and Salimans, 2022) is a standard technique in diffusion models for enhancing conditional generation. The models are trained with randomly dropped conditions to mix conditional and unconditional data. During inference, CFG is applied by linearly combining predictions from the label-conditional and unconditional models to enhance generation quality.
@@
$\mathcal{L}_{\text{SCM}}^{g}({{\theta}})=\mathbb{E}_{t,l,s,{x}_{t},c}\left[\frac{w_{\text{SCM}}(t,l,s,\text{MSE})}{n}\left\|{f}_{{{\theta}}}({x}_{t},t,s,c)-{f}_{{{\theta}}^{-}}\bigl({x}_{t}+v_{g}(x_{t},t,c)(l-t),l,s,c\bigr)\right\|_{2}^{2}\right],$ (27)
@@
However, since the unconditional velocity field is generally intractable during conditional training, $v_{g}(x_{t},t,c)$ is not directly accessible. To address this, we train the network to concurrently predict the unconditional velocity field. Specifically, we randomly replace the condition c{c} with an empty label $\phi$ (with a probability of $0.1$) and update the model using the unconditional Flow Matching loss (Equation 17) and the unconditional solution consistency loss (Equation 20). Consequently, we can approximate $v(x_{t},t)$ using the model prediction $v_{\text{uncond}}=\partial_{2}a(t,t){x}_{t}+\partial_{2}b(t,t){F}_{{\theta}^{-}}({x}_{t},t,t,\phi)$.
For data points with a non-empty condition c{c}, we compute the guided losses by substituting the velocity term $v_{g}(x_{t},t,c)$ in Equation 26 and Equation 27 with the estimator $w(\alpha_{t}^{\prime}{x}_{0}+\beta_{t}^{\prime}{x}_{1})+(1-w){v}_{\text{uncond}}$, where $w$ is the CFG strength. Here, $\alpha_{t}^{\prime}{x}_{0}+\beta_{t}^{\prime}{x}_{1}$ is used to replace the intractable conditional marginal velocity field $v(x_{t},t\mid c)$, similar to the unconditional formulation.
Notably, the term $w(\alpha_{t}^{\prime}{x}_{0}+\beta_{t}^{\prime}{x}_{1})+(1-w){v}_{\text{uncond}}$ typically exhibits higher variance than the original term $\alpha_{t}^{\prime}{x}_{0}+\beta_{t}^{\prime}{x}_{1}$, primarily due to the scaling effect of $w$. This increased variance can hinder training convergence and degrade final performance. We observe that the model learns the guided velocity field via Equation 27 when conditioned on c{c}. Therefore, we can obtain a model-predicted guided velocity ${v}_{\text{guided}}$ via $\partial_{2}a(t,t){x}_{t}+\partial_{2}b(t,t){F}_{{\theta}^{-}}({x}_{t},t,t,{c})$, and then employ
@@
The inference process of our model is straightforward. Since ${f}_{{\theta}}({x}_{t},t,s)\!=\!a(t,s){x}_{t}+b(t,s){F}_{{\theta}}({x}_{t},t,s)$ is a neural solution function to the velocity ODE, we only need to sample ${x}_{1}\sim\mathcal{N}({0},{I}_{n})$ and apply this function with $t=1$ and $s=0$ to obtain the clean data in a single step. Furthermore, similar to consistency models (Song et al., 2023), our model supports multi-step sampling by adding noise to the predicted data and recursively applying the solution function.
Section 5 Experiments

### v2

@@ Abstract
Figure 1: Visual samples and quantitative comparison. (a) One-step samples generated by our Solution Flow Models on the ImageNet $\times256$ dataset. (b) With the same Diffusion Transformer (peebles2023scalable) architecture and the same number of training epochs, our models (SoFlow) consistently achieve better 1-NFE FID-50K scores than MeanFlow (geng2025meanflow) models on the ImageNet $\times256$ dataset.
1 Introduction
Diffusion models (sohl2015deep; song2019generative; ho2020denoising; song2020score) and Flow Matching models (lipman2022flow; liu2022flow; albergo2022building) have emerged as foundational frameworks in generative modeling. Diffusion models operate by systematically adding noise to data and then learning a denoising process to generate high-quality samples. Flow Matching offers a more direct alternative, modeling the velocity fields that transport a simple prior distribution to a complex data distribution. Despite their power and success across various generative tasks (esser2024scaling; ma2024sit; polyak2024movie), both approaches rely on an iterative, multi-step sampling process, which hinders their generation efficiency.
Addressing this latency is becoming a key area of research. Consistency Models (song2023consistency; song2023improved; geng2024consistency; lu2024simplifying) and related techniques (kim2023consistency; wang2025stable; frans2024one; heek2024multistep) have gained prominence by enabling rapid, few-step generation. These methods learn a direct mapping from any point on a generative trajectory to a consistent, “clean” output, bypassing the need for iterative refinement. However, this paradigm introduces significant challenges. Consistency models trained from scratch often fail to leverage Classifier-Free Guidance (CFG) for enhancing sample quality, and they are further hindered by instability caused by changing optimization targets. Recent works (peng2025flow; geng2025meanflow) address this instability by incorporating a Flow Matching loss. However, this approach introduces a new computational bottleneck: it relies on Jacobian-vector product (JVP) calculations, which are much less optimized than forward propagation in frameworks such as PyTorch.
@@ 1 Introduction
In this work, we introduce a new approach for one-step generation that avoids these limitations. Instead of relying on iterative ODE solvers, we propose directly learning the solution function of the velocity ODE defined by Flow Matching to generate high-quality samples, as shown in Figure 1, LABEL:fig:imagenet_post, and LABEL:fig:cifar_post. We denote this function as $f(x_{t},t,s)$, which explicitly maps a state $x_{t}$ at time $t$ to its evolved state $x_{s}$ at another time $s$. To learn it with a parameterized model $f_{\theta}(x_{t},t,s)$, we first analyze the properties that enable a neural network to serve as a valid solution function.
@@ 2 Related Work
Diffusion models (sohl2015deep; song2020score; kingma2021variational; karras2022elucidating) and Flow Matching (lipman2022flow; liu2022flow) are widely adopted generative modeling frameworks. These approaches either progressively corrupt data with noise and train a neural network to denoise it, or learn a velocity field that governs a transformation from the data distribution to a simple prior. They have been scaled successfully for image generation (rombach2022high; saharia2022photorealistic; podell2023sdxl) and video generation (ho2022video; brooks2024video) tasks. Stochastic interpolants (albergo2022building; albergo2023stochastic) build upon these concepts by explicitly defining stochastic trajectories between the data and prior distributions and aligning their associated velocity fields to enable effective distributional transport.
@@
Reducing the number of sampling steps has become an active research direction, driven by the need for faster generation and better theoretical insights. Prior efforts fall into two main streams: (i) distillation-based approaches that compress pre-trained multi-step models into few-step generators (salimans2022progressive; sauer2024adversarial; geng2023one; luo2023diff; yin2024one; zhou2024score); and (ii) from-scratch training, which learns fast samplers without teachers, most prominently through the consistency-training family (CMs, iCT, ECT, sCT) (song2023consistency; song2023improved; geng2024consistency; lu2024simplifying).
Consistency Models (CMs) collapse noised inputs directly to clean data, enabling one- or few-step generation (song2023consistency). While first applied mainly in distillation (luo2023latent; geng2024consistency), later works established that they can also be trained from scratch via consistency training (song2023improved). Building on this, iCT (song2023improved) simplifies objectives and improves stability with robust losses and teacher-free training, while ECT (geng2024consistency) introduces a continuous-time ODE formulation that unifies CMs and diffusion. The recent sCT framework (lu2024simplifying) further stabilizes continuous-time consistency training, scaling CMs up to billion-parameter regimes. Beyond fixed start and end settings, newer approaches extend CMs to arbitrary timestep transitions, aligning better with continuous-time dynamics.
Specifically, Shortcut Models (frans2024one) condition on noise level and step size to flexibly support both one- and few-step sampling with shared weights. MeanFlow (geng2025meanflow) introduces interval-averaged velocities and analyzes the relationship between averaged and instantaneous velocities. IMM (zhou2025inductive) matches moments across transitions in a single-stage objective, reducing variance and avoiding multi-stage distillation. Flow-Anchored CMs (peng2025flow) regularize shortcut learning with a Flow Matching anchor, improving stability and generalization. On the distillation side, Align Your Flow (sabour2025align) unifies CM and FM into continuous-time flow maps effective across arbitrary step counts, further enhanced with autoguidance and adversarial fine-tuning. Finally, Transition Models (wang2025transition) reformulate generation around exact finite-interval dynamics, unifying few- and many-step regimes and achieving monotonic quality gains as step budgets increase.
3 Preliminary: Flow Matching
@@
However, directly optimizing a neural network with this loss is impractical since the real velocity field ${v}({x}_{t},t)$ is a conditional expectation given ${x}_{t}$. To overcome this challenge, a conditional variant of the Flow Matching loss is introduced (lipman2022flow; liu2022flow):
@@
4 Solution Flow Models
@@ 4.1 Formulations
The expression is equal to 0{0} due to the conditions in Eq. 6 and Eq. 9. Thus, we know that the model ${f}_{{{\theta}}}({x}_{t},t,s)$ is indeed the true solution function ${f}({x}_{t},t,s)$ by:
@@ 4.2 Learning Objectives
We now consider how to construct efficient objectives for neural networks to learn the ground truth solution function, according to the two conditions mentioned in the previous section. To ensure the first boundary condition in Eq. 8 is satisfied, we adopt the following parameterization:
@@
where $a(t,s)$ and $b(t,s)$ are continuously differentiable scalar functions satisfying $a(t,t)=1$ and $b(t,t)=0$ for all $t\in[0,1]$, and ${F}_{{{\theta}}}({x}_{t},t,s)$ denotes a raw neural network. Following the notation in the last section, we use $\partial_{1}a(t,s),\partial_{2}a(t,s),\partial_{1}b(t,s),\partial_{2}b(t,s)$ to denote the partial derivatives of $a(t,s)$ and $b(t,s)$ with respect to the first and second variables. In our experiments, we choose two specific parameterizations, including the Euler parameterization and the trigonometric parameterization:
@@
Trivially, these two parameterizations satisfy the boundary condition ${f}_{{\theta}}({x}_{t},t,t)={x}_{t}$. Then, we construct two loss functions for training using Eq. 9.
@@
Firstly, we consider a special case of Eq. 9 when $t=s$. Recall that the boundary condition Eq. 8 gives ${f}_{{\theta}}({x}_{t},t,t)={x}_{t},\forall{x}_{t}\in\mathbb{R}^{n},t\in[0,1]$. Thus, we have:
@@
where ${I}_{n}\in\mathbb{R}^{n\times n}$ is the identity matrix. Then Eq. 9 can be simplified to ${v}({x}_{t},t)=\partial_{3}{f}_{{\theta}}({x}_{t},t,t)$. Under our parameterization mentioned above, we have
@@
Previous works (geng2024consistency; geng2025meanflow) have demonstrated that choosing an adaptive weighting function is beneficial for few-step generative models. Following their approach, we choose
@@
The original Flow Matching framework samples $t$ uniformly, while more recent works (esser2024scaling; geng2025meanflow) suggest sampling $t$ from a logit-normal distribution. We also sample $t$ from $\sigma(\mathcal{N}(\mu_{\text{FM}},\sigma_{\text{FM}}^{2}))$, where $\sigma(\cdot)$ and $\mathcal{N}$ represent the sigmoid function and normal distribution.
@@
We now consider how to build a training target for the $s<t$ situation using Eq. 9. According to the Taylor expansion, we have an approximation equation:
@@
$\mathcal{L}_{\text{SCM}}({{\theta}})\!=\!\mathop{\mathbb{E}}_{\begin{subarray}{c}t,l,s,\\ {x}_{0},{x}_{1},{x}_{t}\end{subarray}}\left[\frac{w_{\text{SCM}}(t,l,s,\text{MSE})}{n}\left\|{f}_{{{\theta}}}({x}_{t},t,s)-{f}_{{{\theta}}^{-}}\bigl({x}_{t}+(\alpha_{t}^{\prime}{x}_{0}+\beta_{t}^{\prime}{x}_{1})(l-t),l,s\bigr)\right\|_{2}^{2}\right],$ (20)
@@
where $k,K$ represent the current and total training steps, respectively. The function $r(k,K)$ represents a monotonically decreasing schedule that gradually moves $l$ towards $t$ throughout training. To avoid numerical issues, $l$ is clamped to ensure $l<t-10^{-4}$. This schedule decreases from an initial value $r_{\text{init}}$ to an end value $r_{\text{end}}$. In our ablation studies, we test exponential, cosine, linear, and constant schedules. For more implementation details, please refer to LABEL:appa.
@@ 4.3 Classifier-Free Guidance
Classifier-Free Guidance (CFG) (ho2022classifier) is a standard technique in diffusion models for enhancing conditional generation. The models are trained with randomly dropped conditions to mix conditional and unconditional data. During inference, CFG is applied by linearly combining predictions from the label-conditional and unconditional models to enhance generation quality.
@@
$\mathcal{L}_{\text{SCM}}^{g}({{\theta}})=\mathop{\mathbb{E}}_{t,l,s,{x}_{t},c}\left[\frac{w_{\text{SCM}}(t,l,s,\text{MSE})}{n}\left\|{f}_{{{\theta}}}({x}_{t},t,s,c)-{f}_{{{\theta}}^{-}}\bigl({x}_{t}+v_{g}(x_{t},t,c)(l-t),l,s,c\bigr)\right\|_{2}^{2}\right],$ (27)
@@
However, since the unconditional velocity field is generally intractable during conditional training, $v_{g}(x_{t},t,c)$ is not directly accessible. To address this, we train the network to concurrently predict the unconditional velocity field. Specifically, we randomly replace the condition c{c} with an empty label $\phi$ (with a probability of $0.1$) and update the model using the unconditional Flow Matching loss (Eq. 17) and the unconditional solution consistency loss (Eq. 20). Consequently, we can approximate $v(x_{t},t)$ using the model prediction $v_{\text{uncond}}=\partial_{2}a(t,t){x}_{t}+\partial_{2}b(t,t){F}_{{\theta}^{-}}({x}_{t},t,t,\phi)$.
For data points with a non-empty condition c{c}, we compute the guided losses by substituting the velocity term $v_{g}(x_{t},t,c)$ in Eq. 26 and Eq. 27 with the estimator $w(\alpha_{t}^{\prime}{x}_{0}+\beta_{t}^{\prime}{x}_{1})+(1-w){v}_{\text{uncond}}$, where $w$ is the CFG strength. Here, $\alpha_{t}^{\prime}{x}_{0}+\beta_{t}^{\prime}{x}_{1}$ is used to replace the intractable conditional marginal velocity field $v(x_{t},t\mid c)$, similar to the unconditional formulation.
Notably, the term $w(\alpha_{t}^{\prime}{x}_{0}+\beta_{t}^{\prime}{x}_{1})+(1-w){v}_{\text{uncond}}$ typically exhibits higher variance than the original term $\alpha_{t}^{\prime}{x}_{0}+\beta_{t}^{\prime}{x}_{1}$, primarily due to the scaling effect of $w$. This increased variance can hinder training convergence and degrade final performance. We observe that the model learns the guided velocity field via Eq. 27 when conditioned on c{c}. Therefore, we can obtain a model-predicted guided velocity ${v}_{\text{guided}}$ via $\partial_{2}a(t,t){x}_{t}+\partial_{2}b(t,t){F}_{{\theta}^{-}}({x}_{t},t,t,{c})$, and then employ
@@
The inference process of our model is straightforward. Since ${f}_{{\theta}}({x}_{t},t,s)=a(t,s){x}_{t}+\allowbreak b(t,s){F}_{{\theta}}({x}_{t},t,s)$ is a neural solution function to the velocity ODE, we only need to sample ${x}_{1}\sim\mathcal{N}({0},{I}_{n})$ and apply this function with $t=1$ and $s=0$ to obtain the clean data in a single step. Furthermore, similar to consistency models (song2023consistency), our model supports multi-step sampling by adding noise to the predicted data and recursively applying the solution function.
5 Experiments